# Generative Artificial Intelligence-Supported Pentesting: A Comparison between Claude Opus, GPT-4, and Copilot

### v1

@@ Abstract
The advent of Generative Artificial Intelligence (GenAI) has brought a significant change to our society. GenAI can be applied across numerous fields, with particular relevance in cybersecurity. Among the various areas of application, its use in penetration testing (pentesting) or ethical hacking processes is of special interest. In this paper, we have analyzed the potential of leading generic-purpose GenAI tools—Claude Opus, GPT-4 from ChatGPT, and Copilot—in augmenting the penetration testing process as defined by the Penetration Testing Execution Standard (PTES). Our analysis involved evaluating each tool across all PTES phases within a controlled virtualized environment. The findings reveal that, while these tools cannot fully automate the pentesting process, they provide substantial support by enhancing efficiency and effectiveness in specific tasks. Notably, all tools demonstrated utility; however, Claude Opus consistently outperformed the others in our experimental scenarios.
@@ 1 Introduction
The advent of Large Language Models (LLMs) and their implementation in systems such as ChatGPT (OpenAI) has brought about a global revolution, reshaping various aspects of society [1, 2]. Alongside the rise of ChatGPT, other LLM-based conversational systems have also emerged [3], including Copilot (Microsoft), Google Gemini, LlamaChat (Meta) and Claude Opus.
@@
The remainder of this paper is organized as follows: Section 2 provides a brief overview of the PTES methodology. Section 3 reviews related work in the field. In Section 4, we present an overview of the main GenAI tools applicable to the pentesting process. Section 5 details our analysis of the use of Claude Opus, GPT-4 from ChatGPT, and Copilot across the different phases of PTES. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 3 Related Work
The work of Hilario et al. [21] is the first detailed study to explore the use of Generative AI (GenAI), specifically ChatGPT, for pentesting purposes. Using a scenario involving Kali Linux and a vulnerable machine named PumpkinFestival from VulnHub, they executed the various steps of the penetration testing process, from reconnaissance to exploitation, utilizing Shell_GPT (sgpt) [^3] https://github.com/TheR1D/shell_gpt. Their tests across different phases demonstrate that ChatGPT can efficiently assist in the pentesting process.
@@ 5 Pentesting using GenAI tools
In this section, we present a comparison of various generic-purpose Generative AI (GenAI) tools to analyze how they support pentesters during the pentesting process within a defined scenario. We first describe the scenario where the tools were tested, followed by the methodology employed, and conclude with the results obtained from the pentesting process.
@@ 5.1 Testing environment: GOAD
The testing environment used for this study is the GOAD (Goad of Active Directory) [^7] https://github.com/Orange-Cyberdefense/GOAD/tree/main, a large-scale laboratory designed to simulate environments with numerous vulnerabilities related to Windows Server systems. This setup consisted of five virtual machines, two forests, three domains, and numerous user accounts, as illustrated in Figure 1.
@@ 5.3 Comparison of GenAI tools in PTES phases
To establish a baseline, the following table summarizes key aspects such as cost, token usage, and knowledge coverage:
@@ 5.3.1 Reconnaissance
Could you give me more information about the high-ranking positions in this company, as well as their social media accounts? Do they have any outsourced services? Could you find information about their servers, DNS services, or employees? Could you give me information about Alejandro Cano [^8] Alejadro Cano is a Technical Director and partner in Legitec
@@
GPT-4
GPT-4 provides basic information about Legitec, including details previously mentioned. Additionally, it supplements this information with the company’s phone number, email address, areas of expertise, and physical address. Furthermore, it generates a list of employees along with their respective positions.
For the final query, GPT-4 created a concise yet detailed report that included information about experience, education, social media profiles, and some personal data.
@@
The data obtained during the reconnaissance phase is relatively easy to find through social media platforms or traditional search engines like Google. The primary advantage of using generative AI tools for these queries lies in the significant time savings they offer. In this case, where only Legitec’s information was analyzed, both Claude Opus and GPT-4 demonstrated similar capabilities. These tools efficiently gathered details about the generic data of the company. Although Claude Opus provide more details of the company. On the other hand, GPT-4 is able to provide a report on a specific individual, Alejandro Cano, which made it more versatile in this context.
In contrast, Copilot was less effective in gathering relevant information. While it retrieved basic details about Legitec, it did not provide specifics like the company’s location or its specialization in cybersecurity audits. Furthermore, when asked for additional information, such as details about employees or individuals, Copilot refused to respond, rendering it less useful for the reconnaissance phase. While traditional methods of gathering this data, such as through social media or search engines, could be time-consuming, the generative AI tools, particularly GPT-4, offered significant time savings and enhanced efficiency in retrieving valuable information.
@@ 5.3.2 Vulnerability Analysis
Upon completing the vulnerability analysis phase, our evaluation reveals that Claude Opus and GPT-4 are the most effective tools, while Copilot falls behind due to its message limitations. Claude Opus point outs in generating a comprehensive summary of open ports and offers detailed guidance on how to analyze vulnerabilities in these ports. It is particularly effective at identifying a wide range of vulnerabilities compared to the other tools, providing a thorough and actionable analysis. The recommendations provided by Claude Opus allows for a more complete assessment of potential weaknesses in the system (Kerberos, LDAP, SMB, RDP, WinRM, HTTP, and MSQL). Although in some cases is necessary to adjust the commands.
GPT-4 also demonstrates strong capabilities showing potential vulnerabilities. It categorizes the vulnerabilities it identifies into general services and those specific to each host, helping to prioritize actions based on the context. While it is effective in generating actionable insights, it tends to be slightly less comprehensive than Claude Opus in its vulnerability identification, focusing more on categorization rather than providing extensive guidance on specific attack vectors.
@@
In summary, Claude Opus and GPT-4 stand out for their ability to provide detailed and context-specific analysis, with Claude Opus being particularly effective at identifying a broader range of vulnerabilities. Copilot still offers valuable assistance, especially for users looking for multi-functional scripts, but its limitations in data processing and organization reduce its effectiveness in the vulnerability analysis phase. Despite these differences, all three tools share a common strength in their ability to filter and summarize large volumes of data, enabling pentesters to quickly identify critical vulnerabilities in complex systems. To optimize the use of these tools, it is essential to specify the tools and environment being tested to ensure that the recommendations are tailored to the specific context.
@@ 5.3.3 Exploitation
Among the recommendations made by the various GenAI tools that were successfully exploited were Kerberos roasting (an attack aimed at stealing password hashes from Active Directory accounts with a Service Principal Name), AS-REP (Authentication Server Response) Roasting (used to steal password hashes from user accounts with Kerberos pre-authentication disabled, allowing attackers to crack the passwords later without detection), and ADCS (Active Directory Certificate Services).
@@
Upon concluding the exploitation phase, it was confirmed that Claude Opus and GPT-4 proved especially useful for generating specific attacks, offering great precision in formulating them. In contrast, Copilot lagged behind due to its limitations in generating customized commands. Both Claude Opus and Copilot were effective in generating payloads or commands for certain attacks, but Claude Opus stood out for its ability to adapt to the context of the environment. It provided tailored guidance for complex attacks such as Kerberos roasting, AS-REP roasting, and ADCS exploitation, generating highly specific and actionable commands.
The ideal tool for this phase has been Claude Opus, thanks to its precision, ease of use, and its great ability to maintain the context of all conversations. Its capacity to adapt responses to the specific environment and exploit vulnerabilities with greater specificity makes it the most effective tool for exploitation. While GPT-4 also offered useful guidance, it lacked the same level of customization and refinement seen in Claude Opus. Copilot, on the other hand, was less effective in this phase, providing more generic and less adaptable responses, which made it less useful for complex exploitation tasks.
@@ 5.3.4 Post Exploitation
Due to Windows Defender, most of the suggested commands could not be executed. As a result, we requested methods to disable it through PowerShell. While it provided several options, we were unable to deactivate Windows Defender.
@@
GPT-4
The commands generated by GPT-4 were very similar to those in the previous section. As before, the initial commands were adapted for a Windows device, with an approach focused on executing Mimikatz for performing SID Injection. However, the subsequent commands were neither well formulated nor coherent, as they instructed us to execute a non-existent .exe file. The generated command is shown below:
@@
As mentioned earlier, we requested alternative tools to execute from Kali, with recommendations including a Metasploit exploit101010use exploit/windows/smb/ms08_067_netapi, Impacket, Winexe, and CrackMapExec. As a result, the attack once again involved obtaining verification through LDAP signing111111crackmapexec smb 192.168.56.11 -u jon.snow -p iknownothing -d north.sevenkingdoms.local -x ’ipconfig’. After authenticating, we obtained the certificate and used the TGT to access the system.
@@
In the context of post-exploitation, Claude Opus, GPT-4, and Copilot each demonstrated distinct strengths and limitations. Although privilege escalation on the target system was not successfully achieved, the methods applied were logically executed, and the tools were capable of correctly formulating the indicated attacks after two or more messages in the conversation. These tools provided precise commands to execute the mentioned attack, although the post-exploitation phase remains to be studied in more depth, as it was only tested for privilege escalation, leaving out other relevant actions such as hiding performed actions and lateral movements.
@@
GPT-4, while equally capable in generating useful and powerful commands, showed a slight difference compared to Claude Opus in terms of the number of tools recommended and its adaptability. GPT-4 successfully provided well-structured attack vectors. However, it occasionally required manual adjustments, as some commands were more generic compared to Claude Opus, which had the ability to offer more contextually specific suggestions. Despite this, GPT-4 was still a valuable asset in the post-exploitation phase, especially when it came to generating additional tools and attack options.
In contrast, Copilot lagged behind Claude Opus and GPT-4 in terms of customization and adaptability. Copilot provided basic, useful commands, but its lack of context-awareness and inability to generate highly specific commands meant that it was not as effective in complex post-exploitation tasks. While Copilot’s commands were still functional, they required more manual intervention to tailor them to the specific environment. Its responses were less refined compared to Claude Opus and GPT-4, making it the least effective of the three tools for the post-exploitation phase.
In conclusion, while all three tools contributed to the post-exploitation phase, Claude Opus and GPT-4 were clearly the more effective and adaptable tools. Their ability to generate precise, context-aware commands and offer a wider range of attack vectors made them better choices for post-exploitation tasks. Copilot, though useful for basic tasks, was hindered by its lack of customization and context sensitivity, which limited its effectiveness in more complex exploitation scenarios.
@@ 5.3.5 Reporting
GPT-4
GPT-4 produced results similar to those of Claude Opus, beginning with the creation of an executive summary that outlined the identified vulnerabilities, categorizing them by CVE and providing more detailed explanations. For instance, the following text presents the output it generated:
@@
In the reporting phase, all three GenAI tools—Claude Opus, GPT-4, and Copilot—generated executive summaries and identified vulnerabilities, but with varying levels of detail. Claude Opus provided comprehensive summaries with detailed CVE references and useful recommendations, though some were general. GPT-4 offered similar results, with slightly more detailed explanations and categorization of vulnerabilities, making it particularly useful for both reporting and analysis. In contrast, Copilot generated more basic reports, lacking specific CVE references and offering less detailed vulnerability explanations. Its recommendations were more generic and limited in scope, reducing its effectiveness in complex pentesting scenarios.
Overall, Claude Opus and GPT-4 delivered more detailed and actionable insights than Copilot, highlighting differences in the depth of reporting provided by these tools.
@@ 5.4 Global assessment
GPT-4, while similar in usability and adaptability, falls slightly short in areas like memory retention and command customization compared to Claude Opus. Despite this, it remains a strong choice for the Exploitation, Post-Exploitation, and Reporting phases, where it performs well in generating structured commands and offering useful suggestions. While GPT-4 may occasionally recommend commands/scripts that are not available or require manual adjustments, particularly in adapting commands for specific contexts, it is still an extremely useful tool in the post-exploitation and reporting phases, providing clear, actionable insights.
Copilot, on the other hand, was the least comfortable to use due to its limitations in expert mode, where only five messages per conversation are allowed. This was especially problematic when explaining the environment or seeking more specificity, as it frequently provided generic responses. The lack of the ability to revisit previous chats further hindered its usability during complex tasks. Additionally, Copilot falls short in command customization compared to Claude Opus and GPT-4, and its generated commands lacked the specificity needed for complex exploitation or vulnerability identification. Despite these drawbacks, Copilot can still serve as an assistant for basic tasks or when specific attacks, such as those targeting MSSQL, are required.
In Table 2 we sum up the overall strengths and weaknesses we have found during the process.
@@
GPT-4 - Produces well-structured outputs and summaries.
@@
Based on the analysis, the Table 3 summarizes the most suitable GenAI tools for each phase of PTES, along with the specific strengths or weaknesses of each tool. Claude Opus consistently stands out, particularly for its ability to handle general reconnaissance, propose customized commands for exploitation, and offer detailed reporting with actionable recommendations. GPT-4, although occasionally requiring refinements, also proves effective in multiple phases—especially exploitation and reporting—where its structured outputs and adaptability can enhance productivity. Copilot, meanwhile, exhibits notable constraints in both contextual understanding and command customization, limiting its utility to simpler or more narrowly scoped tasks. Despite these differences, all three tools offer time-saving features that can supplement the expertise of human pentesters, underscoring the value of integrating GenAI solutions into the pentesting workflow.
Phase Claude Opus GPT-4 Copilot
@@
In conclusion, Claude Opus emerges as the most versatile and effective tool for penetration testing due to its precision, adaptability, and ability to generate context-specific commands across all phases. GPT-4 is a solid tool for generating structured commands and reports, though it requires some refinement in complex exploitation tasks. Copilot, while useful for basic tasks, struggles with customization and context-awareness, making it the least effective choice for more advanced pentesting scenarios.

### v2

@@ Abstract
The advent of Generative Artificial Intelligence (GenAI) has brought a significant change to our society. GenAI can be applied across numerous fields, with particular relevance in cybersecurity. Among the various areas of application, its use in penetration testing (pentesting) or ethical hacking processes is of special interest. In this paper, we have analyzed the potential of leading generic-purpose GenAI tools—Claude Opus, GPT-4 Turbo from ChatGPT, and Copilot—in augmenting the penetration testing process as defined by the Penetration Testing Execution Standard (PTES). Our analysis involved evaluating each tool across all PTES phases within a controlled virtualized environment. The findings reveal that, while these tools cannot fully automate the pentesting process, they provide substantial support by enhancing efficiency and effectiveness in specific tasks. Notably, all tools demonstrated utility; however, Claude Opus consistently outperformed the others in our experimental scenarios.
@@ 1 Introduction
The advent of Large Language Models (LLMs) and their implementation in systems such as ChatGPT (OpenAI) has brought about a global revolution, reshaping various aspects of society [1, 2]. Alongside the rise of ChatGPT, other LLM-based conversational systems have also emerged [3], including Copilot (Microsoft), Google Gemini, LlamaChat (Meta), and Claude Opus.
@@
The remainder of this paper is organized as follows: Section 2 provides a brief overview of the PTES methodology. Section 3 reviews related work in the field. In Section 4, we present an overview of the main GenAI tools applicable to the pentesting process. Section 5 details our analysis of the use of Claude Opus, GPT-4 Turbo from ChatGPT, and Copilot across the different phases of PTES. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 3 Related Work
The work of Hilario et al. [21] is the first detailed study to explore the use of GenAI, specifically ChatGPT, for pentesting purposes. Using a scenario involving Kali Linux and a vulnerable machine named PumpkinFestival from VulnHub, they executed the various steps of the penetration testing process, from reconnaissance to exploitation, utilizing Shell_GPT (sgpt) [^3] https://github.com/TheR1D/shell_gpt. Their tests across different phases demonstrate that ChatGPT can efficiently assist in the pentesting process.
@@ 5 Pentesting using GenAI tools
In this section, we present a comparison of various generic-purpose GenAI tools to analyze how they support pentesters during the pentesting process within a defined scenario. We first describe the scenario where the tools were tested, followed by the methodology employed, and conclude with the results obtained from the pentesting process.
@@ 5.1 Testing environment: GOAD
The testing environment used for this study is the GOAD (Game of Active Directory) [31], a large-scale laboratory designed to simulate environments with numerous vulnerabilities related to Windows Server systems. This setup consisted of five virtual machines, two forests, three domains, and numerous user accounts, as illustrated in Figure 1.
@@
5.1.1 Environment Configuration Details
The GOAD environment deployed in this study consisted of the following technical specifications:
  • Virtual Machines: Five Windows Server virtual machines.
  • Operating Systems deployed: Windows Server 2019 Standard and Windows Server 2016 Standard.
  • Hardware Specifications per VM: 2 virtual CPUs, 4 GB RAM, 40 GB disk space.
  • Network Configuration: Internal virtual network with isolated segments with the IP ranges: 192.168.56.11-13, 192.168.56.22-23.
  • Tools and Versions Used: Kali Linux Rolling Release (kernel 6.6) as attacking machine, Certipy v4.9.0, Enum4linux-ng v1.2, CrackMapExec v5.4.0, Impacket v0.11.0 (addcomputer.py, GetNPUsers.py, GetUserSPNs.py, getST.py, lookupsid.py, ntlmrelayx.py, psexec.py), NetExec, Nmap v7.94, PetitPotam, and Rubeus.
  • Domains: sevenkingdoms.local, north.sevenkingdoms.local, essos.local.
All the details and the write-up can be found in the GitHub of the project [31].
@@ 5.3 Comparison of GenAI tools in PTES phases
To establish a baseline, Table 1 summarizes key aspects such as cost, token usage, and knowledge coverage.
@@ 5.3.1 Reconnaissance
Could you give me more information about the high-ranking positions in this company, as well as their social media accounts? Do they have any outsourced services? Could you find information about their servers, DNS services, or employees? Could you give me information about Alejandro Cano [^7] Alejadro Cano is a Technical Director and partner in Legitec?
@@
GPT-4 Turbo
GPT-4 Turbo provides basic information about Legitec, including details previously mentioned. Additionally, it supplements this information with the company’s phone number, email address, areas of expertise, and physical address. Furthermore, it generates a list of employees along with their respective positions.
For the final query, GPT-4 Turbo created a concise yet detailed report that included information about experience, education, social media profiles, and some personal data.
@@
The data obtained during the reconnaissance phase is relatively easy to find through social media platforms or traditional search engines like Google. The primary advantage of using GenAI tools for these queries lies in the significant time savings they offer. In this case, where only Legitec’s information was analyzed, both Claude Opus and GPT-4 Turbo demonstrated similar capabilities. These tools efficiently gathered details about the generic data of the company. Although Claude Opus provide more details of the company. On the other hand, GPT-4 Turbo is able to provide a report on a specific individual, Alejandro Cano, which made it more versatile in this context.
In contrast, Copilot was less effective in gathering relevant information. While it retrieved basic details about Legitec, it did not provide specifics like the company’s location or its specialization in cybersecurity audits. Furthermore, when asked for additional information, such as details about employees or individuals, Copilot refused to respond, rendering it less useful for the reconnaissance phase. While traditional methods of gathering this data, such as through social media or search engines, could be time-consuming, the generative AI tools, particularly GPT-4 Turbo, offered significant time savings and enhanced efficiency in retrieving valuable information.
@@ 5.3.2 Vulnerability Analysis
Upon completing the vulnerability analysis phase, our evaluation reveals that Claude Opus and GPT-4 Turbo are the most effective tools, while Copilot falls behind due to its message limitations. Claude Opus point outs in generating a comprehensive summary of open ports and offers detailed guidance on how to analyze vulnerabilities in these ports. It is particularly effective at identifying a wide range of vulnerabilities compared to the other tools, providing a thorough and actionable analysis. The recommendations provided by Claude Opus allows for a more complete assessment of potential weaknesses in the system (Kerberos, LDAP, SMB, RDP, WinRM, HTTP, and MSQL). Although in some cases is necessary to adjust the commands.
GPT-4 Turbo also demonstrates strong capabilities showing potential vulnerabilities. It categorizes the vulnerabilities it identifies into general services and those specific to each host, helping to prioritize actions based on the context. While it is effective in generating actionable insights, it tends to be slightly less comprehensive than Claude Opus in its vulnerability identification, focusing more on categorization rather than providing extensive guidance on specific attack vectors.
@@
In summary, Claude Opus and GPT-4 Turbo stand out for their ability to provide detailed and context-specific analysis, with Claude Opus being particularly effective at identifying a broader range of vulnerabilities. Copilot still offers valuable assistance, especially for users looking for multi-functional scripts, but its limitations in data processing and organization reduce its effectiveness in the vulnerability analysis phase. Despite these differences, all three tools share a common strength in their ability to filter and summarize large volumes of data, enabling pentesters to quickly identify critical vulnerabilities in complex systems. To optimize the use of these tools, it is essential to specify the tools and environment being tested to ensure that the recommendations are tailored to the specific context.
@@ 5.3.3 Exploitation
Among the recommendations made by the various GenAI tools that were successfully exploited were Kerberos roasting (an attack aimed at stealing password hashes from Active Directory accounts with a Service Principal Name - SPN), AS-REP (Authentication Server Response) Roasting (used to steal password hashes from user accounts with Kerberos pre-authentication disabled, allowing attackers to crack the passwords later without detection), and ADCS (Active Directory Certificate Services).
@@
Upon concluding the exploitation phase, it was confirmed that Claude Opus and GPT-4 Turbo proved especially useful for generating specific attacks, offering great precision in formulating them. In contrast, Copilot lagged behind due to its limitations in generating customized commands. Both Claude Opus and Copilot were effective in generating payloads or commands for certain attacks, but Claude Opus stood out for its ability to adapt to the context of the environment. It provided tailored guidance for complex attacks such as Kerberos roasting, AS-REP roasting, and ADCS exploitation, generating highly specific and actionable commands.
The ideal tool for this phase has been Claude Opus, thanks to its precision, ease of use, and its great ability to maintain the context of all conversations. Its capacity to adapt responses to the specific environment and exploit vulnerabilities with greater specificity makes it the most effective tool for exploitation. While GPT-4 Turbo also offered useful guidance, it lacked the same level of customization and refinement seen in Claude Opus. Copilot, on the other hand, was less effective in this phase, providing more generic and less adaptable responses, which made it less useful for complex exploitation tasks.
@@ 5.3.4 Post Exploitation
Due to Windows Defender, most of the suggested commands could not be executed. As a result, we requested methods to disable it through PowerShell. Although it provided several options, we were unable to deactivate Windows Defender.
@@
GPT-4 Turbo
The commands generated by GPT-4 Turbo were very similar to those in the previous section. As before, the initial commands were adapted for a Windows device, with an approach focused on executing Mimikatz for performing SID Injection. However, the subsequent commands were neither well formulated nor coherent, as they instructed us to execute a non-existent .exe file. The generated command is shown below:
@@
As mentioned earlier, we requested alternative tools to execute from Kali, with recommendations including the NetAPI exploit from Metasploit999use exploit/windows/smb/ms08_067_netapi, Impacket, Winexe, and CrackMapExec. As a result, the attack once again involved obtaining verification through LDAP signing101010crackmapexec smb 192.168.56.11 -u jon.snow -p iknownothing -d north.sevenkingdoms.local -x ’ipconfig’. After authenticating, we obtained the certificate and used the TGT to access the system.
@@
In the context of post-exploitation, Claude Opus, GPT-4 Turbo, and Copilot each demonstrated distinct strengths and limitations. Although privilege escalation on the target system was not successfully achieved, the methods applied were logically executed, and the tools were capable of correctly formulating the indicated attacks after two or more messages in the conversation. These tools provided precise commands to execute the mentioned attack, although the post-exploitation phase remains to be studied in more depth, as it was only tested for privilege escalation, leaving out other relevant actions such as hiding performed actions and lateral movements.
@@
GPT-4 Turbo, while equally capable in generating useful and powerful commands, showed a slight difference compared to Claude Opus in terms of the number of tools recommended and its adaptability. GPT-4 Turbo successfully provided well-structured attack vectors. However, it occasionally required manual adjustments, as some commands were more generic compared to Claude Opus, which had the ability to offer more contextually specific suggestions. Despite this, GPT-4 Turbo was still a valuable asset in the post-exploitation phase, especially when it came to generating additional tools and attack options.
In contrast, Copilot lagged behind Claude Opus and GPT-4 Turbo in terms of customization and adaptability. Copilot provided basic, useful commands, but its lack of context-awareness and inability to generate highly specific commands meant that it was not as effective in complex post-exploitation tasks. While Copilot’s commands were still functional, they required more manual intervention to tailor them to the specific environment. Its responses were less refined compared to Claude Opus and GPT-4 Turbo, making it the least effective of the three tools for the post-exploitation phase.
In conclusion, while all three tools contributed to the post-exploitation phase, Claude Opus and GPT-4 Turbo were clearly the more effective and adaptable tools. Their ability to generate precise, context-aware commands and offer a wider range of attack vectors made them better choices for post-exploitation tasks. Copilot, though useful for basic tasks, was hindered by its lack of customization and context sensitivity, which limited its effectiveness in more complex exploitation scenarios.
@@ 5.3.5 Reporting
In this prompt, [Nmap Data] represent the data obtained in the scan made with Nmap. [Enum4 Linux Data] represents the users and groups got with Enum4linux. [Crackmapexec Data] is the SMB data obtained with the tool Crackmapexec. [RPC Users] is the users obtained using the rpcclient tool. Finally, [GetNPUsers Data] are the users hashes and passwords.
@@
GPT-4 Turbo
GPT-4 Turbo produced results similar to those of Claude Opus, beginning with the creation of an executive summary that outlined the identified vulnerabilities, categorizing them by CVE and providing more detailed explanations. For instance, the following text presents the output it generated:
@@
In the reporting phase, all three GenAI tools—Claude Opus, GPT-4 Turbo, and Copilot—generated executive summaries and identified vulnerabilities, but with varying levels of detail. Claude Opus provided comprehensive summaries with detailed CVE references and useful recommendations, though some were general. GPT-4 Turbo offered similar results, with slightly more detailed explanations and categorization of vulnerabilities, making it particularly useful for both reporting and analysis. In contrast, Copilot generated more basic reports, lacking specific CVE references and offering less detailed vulnerability explanations. Its recommendations were more generic and limited in scope, reducing its effectiveness in complex pentesting scenarios.
Overall, Claude Opus and GPT-4 Turbo delivered more detailed and actionable insights than Copilot, highlighting differences in the depth of reporting provided by these tools.
@@ 5.4 Global assessment
GPT-4 Turbo, while similar in usability and adaptability, falls slightly short in areas like memory retention and command customization compared to Claude Opus. Despite this, it remains a strong choice for the Exploitation, Post-Exploitation, and Reporting phases, where it performs well in generating structured commands and offering useful suggestions. While GPT-4 Turbo may occasionally recommend commands/scripts that are not available or require manual adjustments, particularly in adapting commands for specific contexts, it is still an extremely useful tool in the post-exploitation and reporting phases, providing clear, actionable insights.
Copilot, on the other hand, was the least comfortable to use due to its limitations in expert mode, where only five messages per conversation are allowed. This was especially problematic when explaining the environment or seeking more specificity, as it frequently provided generic responses. The lack of the ability to revisit previous chats further hindered its usability during complex tasks. Additionally, Copilot falls short in command customization compared to Claude Opus and GPT-4 Turbo, and its generated commands lacked the specificity needed for complex exploitation or vulnerability identification. Despite these drawbacks, Copilot can still serve as an assistant for basic tasks or when specific attacks, such as those targeting MSSQL, are required.
In Table 2 we sum up the overall strengths and weaknesses we have found during the process. Based on the analysis, Table 3 summarizes the most suitable GenAI tools for each phase of PTES, along with the specific strengths or weaknesses of each tool. Claude Opus consistently stands out, particularly for its ability to handle general reconnaissance, propose customized commands for exploitation, and offer detailed reporting with actionable recommendations. GPT-4 Turbo, although occasionally requiring refinements, also proves effective in multiple phases—especially exploitation and reporting—where its structured outputs and adaptability can enhance productivity. Copilot, meanwhile, exhibits notable constraints in both contextual understanding and command customization, limiting its utility to simpler or more narrowly scoped tasks. Despite these differences, all three tools offer time-saving features that can supplement the expertise of human pentesters, underscoring the value of integrating GenAI solutions into the pentesting workflow.
@@
GPT-4 Turbo - Produces well-structured outputs and summaries.
@@
Phase Claude Opus GPT-4 Turbo Copilot
@@
In conclusion, Claude Opus emerges as the most versatile and effective tool for penetration testing due to its precision, adaptability, and ability to generate context-specific commands across all phases. GPT-4 Turbo is a solid tool for generating structured commands and reports, though it requires some refinement in complex exploitation tasks. Copilot, while useful for basic tasks, struggles with customization and context-awareness, making it the least effective choice for more advanced pentesting scenarios.
5.5 Legal and Ethical Considerations
It is important to point out that this assessment has been made in a simulated environment. However, the usage of GenAI tools in real pentesting assessments raise significant ethical and legal challenges since they introduce some risks such as the GenAI tool could suggest commands that, if misused, could enable unauthorized system access; data privacy concerns due to that sharing sensitive system data with cloud-based LLM services may violate confidentiality agreements or privacy regulations such as GDPR; the AI-generated outputs can include incorrect or non-functional commands, requiring human validation to avoid unintended consequences.
Therefore, it is important that when pentesters use these tools that into account several issues such as validating all AI-generated commands before execution, avoiding inputting sensitive client data into external GenAI tools unless contractual and regulatory conditions allow, maintaining human oversight and critical judgment throughout testing engagements, and, documenting the use of AI in reports to provide transparency regarding methodologies and tools used. These recommendations probably will evolve as this technology matures. Meanwhile, additional research should be made to establish clearer guidelines for responsible AI adoption in cybersecurity.